OpenStudy (cieloveswork):

4x=-26 a. 22 b. -22 c. 6 1/2 d. -6 1/2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (amruta1202):

Option d

OpenStudy (skullpatrol):

hint: use the idea of balance to solve the equation

OpenStudy (amruta1202):

X= -26/4 = -13/2 If represented in mixed fractions it is -6 1/2

OpenStudy (skullpatrol):